1.y(3x^3-x+y)dx+x^2(1-x^2)dy=0
2.y(2x+y^2)dx+x(y^2-x)dy=0
1
Expert's answer
2021-07-19T15:38:21-0400

Solution.

1.

"y(3x^3-x+y)dx+x^2(1-x^2)dy=0"

This is the first order nonlinear ordinary differential equation.

Divide on "x^2(1-x^2)dx."


"y'=\frac{3x^3y-xy+y^2}{x^2(x^2-1)},""y'=\frac{y^2}{x^2(x-1)(x+1)}+\frac{3xy}{(x-1)(x+1)}-\frac{y}{x(x-1)(x+1)},"

"y'-\frac{3x^2-1}{x^3-x}y=\frac{y^2}{x^4-x^2}."

We have Bernuli's equation for n=2.

Divide by "y^2."


"\frac{y'}{y^2}-\frac{3x^2-1}{(x^3-x)y}=\frac{1}{x^4-x^2}."

Replacement "u=\frac{1}{y}," then "u'=-\frac{y'}{y^2}, y'=-u'y^2."


"-\frac{u(3x^2-1)}{x^3-x}-u'=\frac{1}{x^4-x^2},"

"u'+(\frac{3x}{(x-1)(x+1)}-\frac{1}{x(x-1)(x+1)})u=-\frac{1}{x^4-x^2}"

This is the first order linear equation.

Let be "P(x)=\frac{3x}{(x-1)(x+1)}-\frac{1}{x(x-1)(x+1)}."

"\int P(x)dx=\int \frac{3x}{(x-1)(x+1)}-\frac{1}{x(x-1)(x+1)}dx=\ln (x^3-x)+C,"

where C is some constant.

Solve

"u'+P(x)u=0,"

"\frac{du}{u}=-P(x)dx,"

"\int\frac{du}{u}=-\int P(x)dx,"

"\ln u =\ln (x^3-x)+C,"

"u=\frac{C}{x^3-x}."

The solution of "u'+(\frac{3x}{(x-1)(x+1)}-\frac{1}{x(x-1)(x+1)})u=-\frac{1}{x^4-x^2}" find by the method of variation constant in the form "u=\frac{C(x)}{x^3-x}."


"C(x)=\int(-\frac{x^3-x}{x^4-x^2})dx=-\ln x+C."

Thus "u=\frac{-\ln x+C}{x^3-x}."

From here "y=\frac{x-x^3}{\ln x+C}."

Answer. "y=\frac{x-x^3}{\ln x+C}."

2.


"x(y^2-x)dy=y(y^2+2x)dx=0."

This is the first order nonlinear ordinary differential equation.

Replacement "y=\sqrt{z}," then "dy=\frac{dz}{2\sqrt{z}}." We will have


"(\frac{c\sqrt{z}}{2}-\frac{x^3}{2\sqrt z})dz=(-z^{\frac{3}{2}}-2x\sqrt z)dx."

Replacement "u=\frac{z}{x}," then

"x=ux, dz=udx+xdu."

We will have


"(\frac{\sqrt {u}}{2}-\frac{1}{2\sqrt u}x^{3/2}(udx+xdu)=(-u^{3/2}-2\sqrt u)x^{3/2}dx,"

"(\frac{\sqrt {u}x^{5/2}}{2}-\frac{x^{5/2}}{2\sqrt u})du=(-\frac{3u^{3/2}x^{3/2}}{2}-\frac{3\sqrt u u^{3/2}}{2})dx."

Divide by "x^{5\/2}" and "-\frac{3u^{3/2}x^{3/2}}{2}-\frac{3\sqrt u u^{3/2}}{2}."

We will have


"(\frac{1}{3u(u+1)}-\frac{1}{3(u+1)})du=\frac{dx}{x}."

"\int(\frac{1}{3u(u+1)}-\frac{1}{3(u+1)})du=\int\frac{dx}{x}."

"\frac{\ln u}{3}-\frac{2\ln (u+1)}{3}=\ln x+C,"

where C is some constant.

From here


"\frac{u}{(u+1)^2}=Cx^3."

Thus


"\frac{z}{(z+x)^2}=Cx^2."

So we have answer


"y^2=Cx^2(y^2+x)^2."

Answer. "y^2=Cx^2(y^2+x)^2."